OpenStudy (tvinsant):

Please help and I will fan and medal... Determine which consecutive integers are the real zeros of each functionary located. /Users/jaylee/Desktop/Screen Shot 2015-11-09 at 1.16.18 PM.png

OpenStudy (tvinsant):

1. f(x)=3x^3-10x^2+22x-4 2. f(x)=x^3-12x^2+17x-9 3. f(x)=x^3-9

OpenStudy (tvinsant):

of each function are located*